OpenStudy (anonymous):

find the oblique asymptotes of, f(x)=x^(2)-5x-7/ x-4

OpenStudy (anonymous):

ummmm............what is oblique??

OpenStudy (anonymous):

what it means is to divide. you will get a quotient and a remainder

OpenStudy (anonymous):

the quotient is your oblique asymptote

OpenStudy (anonymous):

\[\frac{x^2-5x-7}{x-4}=x-1-\frac{11}{x-4}\] and your oblique asymptote is \[y=x-1\]